# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 4.5 Example: Molecular gases
To explain these models, consider a molecular gas immersed in a background fluid. The motion of the Brownian particles can be described by Eq. (77), with $\mathcal{J}[\vec{v},\vec{w}|f(t)]$ a Boltzmann-Enskog (elastic) collision term. The fluid now produces both thermal fluctuations and a nonlinear viscous force [139, 140] $\vec{F}(\vec{v})=\vec{v}\zeta(v)$. The simplest nonlinear model can be obtained by assuming a quadratic dependence of the drag force on the velocity,
@@
Another interesting scenario is that of a binary mixture with linear viscous drag [145]. The coupled Boltzmann-Fokker-Planck equations for both species (again assuming homogeneous states) are
@@
where $\gamma_{i}$ is the friction coefficient and $T_{\rm ext}$ is the background temperature. The presence of the external bath couples the time evolution of the total and partial temperatures of each component, allowing the appearance of the Mpemba phenomenon, even when the initial temperature differences are of the same order of the temperatures themselves.
@@ 4.6 Discussion
Let us clarify some crucial differences between the Markovian and kinetic Mpemba effects. The most important is that all the gas models in the kinetic theory frameworks characterize a driven system that is intrinsically nonequilibrium. Again, the driven-gas framework offers an approach where the system of interest can be treated analytically and is well understood in the thermodynamic limit. Anomalous relaxation effects are therefore observed when driving the system from one nonequilibrium steady state to another, as opposed to the case described in Sec. 2, where the system goes from one Boltzmann equilibrium at some initial temperature $T_{0}$ to another Boltzmann equilibrium at some final temperature $T_{b}$. Also that is in contrast with the Markovian case where the system must be forced to sustain a NESS (See Sec. 8.2.1 for an example) whereas in the granular case the system does not obey an equilibrium Boltzmann distribution.
Figure 26: a) Measures comparison in molecular gases with nonlinear drag: (left) Mpemba occurrence in the temperature evolution and (right) absence of the Mpemba in the relative entropy measure for the same system. Lines are theoretical predictions and open symbols DSMC and filled symbols molecular dynamics simulations. b) Measures comparison in driven Maxwell model of granular gases (left) absence in the total-energy time evolution and (right) presence in the Euclidean distance, defined in the main text, for the same system. Panel a) is reproduced from [138] and panel b) from [156].
Another difference is that, here, the Mpemba effect was defined based on the relaxation of the granular temperature: a crossing in the temperature implies the existence of the “thermal” Mpemba effect. However, one can use other ways to define the effect in granular gases. For example, any relative entropy, such as in the case of the Markovian Mpemba effect, can be used to characterize the “distance” from the final steady state and define a corresponding Mpemba effect. This type of effect is referred to as the “entropic” Mpemba effect in the granular-gas community  [156]. As in the Markovian case [60], for granular gases, the different definitions do not always correspond to the same effects. This was shown in [156], where the two models presented above, the molecular gas with nonlinear drag and anisotropically driven granular models, were considered. It was shown that the “entropic” and “thermal” Mpemba effect are not equivalent in these cases.
@@ 5 Statistics of anomalous thermal relaxations
5.1 From a concrete model to an ensemble of models
@@
The parity of the Mpemba index can be deduced from probability distribution at a few temperatures (see Eqs. (41) and (42)), whereas calculating the Mpemba index itself requires scanning over all temperatures. This is an important advantage in a numerical study, although this estimator is only a lower bound on the probability of having a strong Mpemba effect.
@@ 5.2 How likely is the Mpemba effect in the random energy model?
where the variance, $N\sigma_{E}^{2}$, ensures that the energy and other thermodynamic potentials are extensive. The above choice is the most common; however, other distributions have been studied as well.
The equilibrium distribution for a specific instance of the REM $\{E_{i}\}$ is given by the corresponding Boltzmann distribution
$\pi_{i}(T,\{E_{i}\})=\frac{1}{Z(T,\{E_{i}\})}e^{-E_{i}/T},$ (117)
where $Z(T,\{E_{i}\})=\sum_{i=1}^{L}e^{-E_{i}/T}$ is the partition function of the specific random sample of energy barriers, $\{E_{i}\}$. Each set of energy barriers has its corresponding Helmholtz free energy
@@
Note that here both the partition function and the Helmholtz free energy are random variables, on account of the random choice of the energies. The (quenched) average Helmholtz free energy is given by
@@
where the subscript next to the ensemble average $\langle...\rangle_{\rho_{E}}$ denotes an average over the distribution of energies. Much of the interest in the REM stems from its transition from a delocalized occupancy at high temperatures (where the probability is spread over most states) to a “condensation” of the system in several states for temperatures below the critical temperature, $T_{\rm cr}$[157]. This “freezing” or “condensation” phenomenon is a phase transition, since the average free energy $\langle F(T,\{E_{i}\})\rangle_{E}$ is nonanalytic at the critical temperature [159]. For Gaussian energy statistics, as in  Eq. (116), the critical temperature is $T_{\rm cr}=\sigma_{E}/\sqrt{2\ln 2}$; see  [157].
@@ 5.2.1 The 1D chains with Glauber/Metropolis dynamic ensemble
$W_{i+1,\,i}=\frac{\Gamma}{1+e^{(E_{i+1}-E_{i})/T_{b}}}\quad\text{and}\quad W_{i-1,\,i}=\frac{\Gamma}{1+e^{(E_{i}-E_{i-1})/T_{b}}},$ (120)
@@
$W_{i+1,\,i}=\Gamma\min\left(1,e^{-\frac{E_{i+1}-E_{i}}{T_{b}}}\right)\quad\text{and}\quad W_{i-1,\,i}=\Gamma\min\left(1,e^{-\frac{E_{i-1}-E_{i}}{T_{b}}}\right).$ (121)
@@
The ensemble that we have introduced can be easily studied numerically, and some limiting cases can be studied analytically. In Fig. 27, we numerically evaluate the probabilities for an even direct and inverse Mpemba index as a function of the temperature for the Metropolis dynamics. Repeating the calculation for Glauber dynamics gives indistinguishable results. In both cases, the probability to have a direct Mpemba effect exceeds $0.35$ at high temperatures.
@@ 5.2.2 The random barriers ensemble
$W_{i\!j}(T_{b})=\Gamma e^{-(B_{i\!j}-E_{j})/T_{b}},$ (122)
where $\Gamma^{-1}$ has units of time, $E_{i}$ is the energy of site $i$, $\{B_{i\!j}\}$ parameterizes the barrier between sites $i$ and $j$ with the property ($B_{i\!j}=B_{ji}$), and $T_{b}$ is the bath temperature.
For a given system, the energies $\{E_{i}\}$ are a property of the system itself, regardless of how it is coupled to the environment or the exact nature of the environment. In contrast, the barriers, parameterized by $\{B_{i\!j}\}$, are not just internal properties of the system but rather are functions of the particularities of the environment and its precise coupling to the system. These barriers are often not known or can only be roughly estimated. In the random barrier ensemble, we treat both the energies and the barriers on equal footing, with the barriers also randomly drawn from a distribution,
@@
where $\theta$ is the Heaviside step function. Note that to visualize the set of energies and barriers as a landscape of valleys and peaks, one would need to make sure that $B_{i\!j}\geq\{E_{i},E_{j}\}$, and while the above choice does not facilitate that, without loss of generality one can shift all of the energies by a constant and achieve that indeed $B_{i\!j}$ is higher than $\{E_{i},E_{j}\}$.
Note that, in contrast to the ensembles discussed in Sec. (5.2.1), the system here has “mean-field coupling,” in the sense that state $i$ has non-zero transition rates to all other states. Numerical studies [96] of the probability distribution of the parity of the direct Mpemba effect show that the effect occurs with finite probability both above and below the freezing transition; see Fig. 28.
Figure 28: Lower bound for the direct strong Mpemba effect, showing the likelihood to have an odd number of zeros of $a_{2}$, i.e., for the probability to have a positive Mpemba index parity. The probability is evaluated as a function of the variances of the energy, $\sigma_{E}^{2}\log_{2}L$, and barrier, $\sigma_{B}^{2}\log_{2}L$. The ensemble statistics are shown for two different bath temperatures $k_{B}T_{b}=0.1$ (panel a) and $1.0$ (panel b), for $L=10$ energy levels. Each point was averaged over $2\times 10^{5}$ realizations. The freezing transition temperature is $T_{\rm cr}=\sigma_{E}/\sqrt{2\ln 2}\approx 0.85\,\sigma_{E}$; thus, the figures show that the probability is finite in both phases. The figures are adapted from [5].
@@ 5.2.3 The isotropic ensemble
To evaluate the occurrence of the Mpemba effect across a range of systems, it is essential to determine the eigenvector corresponding to the slowest eigenvalue (the second eigenvector) for each instance in the ensemble. Solving analytically for this second eigenvector as a function of the bath temperature is challenging, particularly when dealing with large matrices. However, a method for approximating the second eigenvector was presented in [5] for a specific ensemble of large random matrices.
To understand the main idea in the isotropic ensemble, let us first consider the two ensembles discussed above in Secs. 5.2.1 and  5.2.2. In both cases, the energies are randomly drawn from a Gaussian distribution. By contrast, the barriers are either fixed by the dynamics (as in the Glauber and Metropolis case) or chosen randomly, too. The eigenvectors of the relaxation dynamics are themselves random because of the randomness of the energies and, in the random-barrier case, also the barriers. In all these cases, finding the slowest eigenvector for large systems and for a given realization at arbitrary temperature can only be done numerically. In the isotropic ensemble, we avoid this problem by directly specifying the statistics of the slowest eigenvector direction rather than the barrier statistics.
To facilitate this process, it is more convenient to work with the symmetrized relaxation matrix: the relaxation rate matrix W can be transformed to a symmetric matrix $\tilde{\textbf{W}}$ via
@@
where $F_{i\!j}(T_{b})=\exp(E_{i}/T_{b})\delta_{i\!j}$. The symmetric matrix $\tilde{\textbf{W}}$ has the same eigenvalues as W and has an orthogonal set of eigenvectors $\vec{f}_{i}=F^{1/2}\vec{v}_{i}$. The eigenvectors $\{\vec{f}_{i}\}$ are orthogonal, $\vec{f}_{i}\cdot\vec{f}_{j}=||\vec{f}_{i}||^{2}\delta_{i\!j}$. Working with a symmetric matrix offers other advantages besides orthogonality; numerical diagonalization algorithms are typically more stable.
The zero-eigenvalue eigenvector of $\tilde{\textbf{W}}$ is given by $(\vec{f}_{1})_{i}\propto e^{-E_{i}/(2T_{b})}$. It is set by choosing the energies $\{E_{i}\}$ and the bath temperature $T_{b}$. The second eigenvector of $\tilde{\textbf{W}}$ is then chosen as $\vec{g}$, a vector of i.i.d. (independent and identically distributed) random variables from a Gaussian distribution, projected onto the space orthogonal to $\vec{f}_{1}$:
@@
To estimate the occurrence of the Mpemba effect, we examine the probability associated with the parity of the Mpemba index. Below, we outline the key steps in estimating the occurrence of the direct Mpemba effect. Similarly, one can calculate the probability to have an inverse Mpemba effect. Both the direct and inverse Mpemba effects are analyzed in Ref. [5].
The parity of the direct Mpemba index, from Eq. (42), is
@@
where $(\vec{u}^{\,\rm dir})_{i}=\exp[E_{i}/(2T_{b})]$ and $w_{i}=\exp[-E_{i}/(2T_{b})]\left(\langle E\rangle_{\vec{\pi}(T_{b})}-E_{i}\right)$, and $\langle E\rangle_{\vec{\pi}(T_{b})}=\sum_{i}E_{i}\,\pi_{i}(T_{b})$ is the average energy with respect to the Boltzmann distribution at $T_{b}$. It is important to note that the vectors $\vec{u}^{\,\rm dir}$ and $\vec{w}$ depend only on the energies and not on $\vec{f}_{2}$. Because of this property, the above formula for the parity has a simple geometric meaning. Decomposing $\vec{f}_{2}$ into $\vec{f}_{||}$ and $\vec{f}_{\perp}$, components that are parallel and perpendicular to $\vec{u}^{\,\rm dir}$, we find
@@
Figure 29: Plane of components of the second eigenvector, $\vec{f}_{2}$, parallel to $\vec{u}^{\,\rm dir}$, $f_{||}$, and perpendicular to $\vec{u}^{\,\rm dir}$, $f_{\perp}$. The direct Mpemba index $\mathcal{I}^{\rm dir}_{\rm M}$ is odd in the double-wedge shaded region of the $f_{||}f_{\perp}-$plane if $\vec{u}^{\,\rm dir}\cdot\vec{w}>0$, and odd in the white regions if $\vec{u}^{\,\rm dir}\cdot\vec{w}<0$. This figure is reproduced from [5].
@@
$\left(\vec{u}^{\,\rm dir}_{\rm iso}\right)_{i}\equiv e^{E_{i}/(2T_{b})}-\frac{Le^{-E_{i}/(2T_{b})}}{Z(T_{b},\{E_{i}\})}.$ (130)
Expressing the vector $\vec{g}$ in terms of two components—one that is parallel to $\vec{u}^{\,\rm dir}_{\rm iso}$, denoted as $g_{||}$ and another that is orthogonal to $\vec{u}^{\,\rm dir}_{\rm iso}$ denoted as $g_{\perp}$—we again geometrically get a double wedge where $\mathcal{P}(\mathcal{I}^{\rm dir}_{\rm M})>0$. Additionally, when estimating the probability of the Mpemba index, it is essential that the entries of $\vec{g}$ be distributed isotropically, and thus also the distribution of entries in its orthogonal projection onto $\vec{f}_{1}$. Since the components $g_{i}$ are independent and identically distributed (i.i.d.) Gaussian random variables, the same applies to $g_{||}$ and $g_{\perp}$. Consequently, we conclude that the area corresponding to the double wedge where $\mathcal{P}(\mathcal{I}^{\rm dir}_{\rm M})>0$ is defined as follows:
@@
Prob($\mathrm{Prob}\left(\mathcal{P}(\mathcal{I}^{\rm dir}_{\rm M})>0\right)\approx\frac{C(\{E_{i}\})}{T_{b}},$ (132)
@@
The numerical results from the random barrier ensemble are compared to the analytical estimates obtained from the isotropic ensemble in Fig. 30. In panel (a), the histograms depict the components of the second eigenvector, $(\vec{f}_{2})_{i}$, generated from the random barrier ensemble in red, alongside the components of the “candidate second eigenvector,” $(\vec{f}^{\,\,\rm{iso}}_{2})_{i}$, shown in green. Both ensembles were created for a fixed energy realization comprising $L=25$ energies drawn from a Gaussian distribution with mean zero and standard deviation $1.5$. Each ensemble consisted of $20,000$ components for the second eigenvector. The barriers were selected from a half-normal distribution with standard deviation of $75.6$. The bath temperature was set to $10$. The components of the random vectors were chosen independently and identically distributed (i.i.d.) from a Gaussian distribution with mean zero and standard deviation $0.19$. As noted in the inset, the two probability distributions exhibit differences at the tails. Figure 30b shows that the expression Eq. (131) obtained for the isotropic ensemble (solid line) captures well the probability of the direct Mpemba index for a random barrier ensemble (points) with a fixed energy realization of $L=25$ energies. This is averaged over $20,000$ randomly selected barrier realizations based on the Eq. (123) statistics. The parameters for the statistics of energies, barriers, and the random vector $\vec{g}$ were the same as those used in Fig. 30a, although the energy sample realization was different. Both panels illustrate that the two ensembles align well when the barrier distribution and the bath temperature exceed the energy spread. In general, more work is needed to establish the range over which the average of the random barrier ensemble closely aligns with the isotropic ensemble.
@@
Figure 30: a) Comparison of the random barrier ensemble and the isotropic ensemble for a quenched energy distribution of $L=25$ energies chosen from a Gaussian distribution with zero mean and standard deviation $\sigma_{E}=1.5$. The bath temperature is $T_{b}=10$. The sizes of the isotropic ensemble are $20,000$ random entries of $\vec{f}^{\,\,\rm iso}_{2}$. The size of the random barrier ensemble is $20,000$ random entries of $\vec{f}_{2}$. The barrier distribution is a half-normal distribution with $\sigma_{B}=75.6$; see Eq. (123). For the isotropic ensemble random vector entries, $g_{i}$ components are i.i.d.s drawn from a Gaussian with zero mean and standard deviation $\sigma_{g}=0.19$. The histogram shows the random barrier ensemble probability of $\vec{f}_{2}$ components, and the green histogram shows the isotropic ensemble histogram of the components for the “candidate second eigenvector,” $\vec{f}^{\,\,\rm iso}_{2}$. The inset emphasizes the tails of the two histograms, where we see the difference between the two ensembles. b) Probability of the direct Mpemba index parity as a function of bath temperature $T_{b}$ for a random barrier ensemble (blue dots) and analytical estimate (orange line). The means and standard deviations of the random vector and the barriers were the same as in a). The quenched sample of $L=25$ energies was different than in a) but obeyed the same statistics (Gaussian ensemble with zero mean and standard deviation 1.5). The two ensembles match well for bath temperature and barrier distributions that are wider than the energy distribution.

### v3

@@ 4.5 Example: Molecular gases
To explain these models, consider a molecular gas immersed in a background fluid. The motion of the Brownian particles can be described by Eq. (77), with $\mathcal{J}[\vec{v},\vec{w}|f(t)]$ a Boltzmann-Enskog (elastic) collision term. The fluid now produces both thermal fluctuations and a nonlinear viscous force [146, 147] $\vec{F}(\vec{v})=-\vec{v}\zeta(v)$. The simplest nonlinear model can be obtained by assuming a quadratic dependence of the drag force on the velocity,
@@
Another interesting scenario is that of a binary mixture with linear viscous drag [152]. The coupled Boltzmann-Fokker-Planck equations for both species, again assuming homogeneous states, are
@@
where $\gamma_{i}$ is the friction coefficient and $T_{\rm ext}$ is the background temperature. The presence of the external bath couples the time evolution of the total and partial temperatures of each component, allowing the appearance of the Mpemba phenomenon, even when the initial temperature differences are of the same order as the temperatures themselves.
@@ 4.6 Discussion
There are some crucial differences between the Markovian and kinetic Mpemba effects. The most important is that all the gas models in the kinetic theory frameworks characterize a driven system that is intrinsically nonequilibrium. Again, the driven-gas framework offers an approach where the system of interest can be treated analytically and is well understood in the thermodynamic limit. Anomalous relaxation effects are therefore observed when driving the system from one nonequilibrium steady state to another, as opposed to the case described in Sec. 2, where the system goes from one Boltzmann equilibrium at some initial temperature $T_{0}$ to another Boltzmann equilibrium at some final temperature $T_{b}$. Note also the contrast with the Markovian case where the system must be forced to sustain a NESS (see Sec. 8.2.1 for an example). In the granular case, the system does not obey an equilibrium Boltzmann distribution.
Figure 26: (a) Measure comparison for molecular gases with nonlinear drag: As we discussed in the text, the relative temperatures of hot and cold states cross (left), but not their relative entropy (right). Lines are theoretical predictions; open symbols are DSMC; and filled symbols are MD. (b) Measure comparison for the case of driven Maxwell model of granular gases. Note that the Mpemba effect is not observed in the total energy for hot and cold states which does not exhibit a crossing (left) whereas the Euclidean distance does (right). Source: Reprinted with permission from [147] and [165] respectively.
Another difference is that, here, the Mpemba effect was defined based on the relaxation of the granular temperature: a crossing in the temperature implies the existence of the “thermal” Mpemba effect. However, we also used other measures to define the Mpemba effect, such as relative entropy (Kullback-Leibler divergence) in Sec. 2.1.4. When using this latter measure, one refers to the “entropic” Mpemba effect in the granular-gas community  [147]. As in the Markovian case [60], for granular gases, the different definitions do not always correspond to the same effects. This was shown in [165], where the two models presented above, the molecular gas with nonlinear drag and anisotropically driven granular models, were considered. It was shown that the “entropic” and “thermal” Mpemba effect are not equivalent in these cases.
@@ 5 Statistics of anomalous thermal relaxations
5.1 From one model to an ensemble
@@
The parity of the Mpemba index can be deduced from probability distribution at a few temperatures (see Eqs. (41) and (42)), whereas calculating the Mpemba index itself requires scanning over all temperatures. This is an important advantage in a numerical study, although this estimator is only a lower bound on the probability to have a strong Mpemba effect.
@@ 5.2 How likely is the Mpemba effect in the Random Energy Model?
where the variance, $N\sigma_{E}^{2}$, ensures that the energy and other thermodynamic potentials are extensive. Although the above choice is the most common, other distributions have been studied as well.
The equilibrium distribution for a specific instance of the REM $\{E_{i}\}$ is given by the corresponding Boltzmann distribution,
$\pi_{i}(T,\{E_{i}\})=\frac{1}{Z(T,\{E_{i}\})}e^{-E_{i}/k_{B}T},$ (117)
where $Z(T,\{E_{i}\})=\sum_{i=1}^{L}e^{-E_{i}/k_{B}T}$ is the partition function of the specific random sample of energy barriers, $\{E_{i}\}$. Each set of energy barriers has its corresponding Helmholtz free energy
@@
Note that the random choice of energies makes both the partition function and the Helmholtz free energy random variables. The (quenched) average Helmholtz free energy is given by
@@
where the ensemble-average subscript $\langle...\rangle_{\rho_{E}}$ denotes an average over the distribution of energies. Much of the interest in the REM stems from its transition from a delocalized occupancy at high temperatures, where the probability is spread over most states, to a “condensation” of the system in several states for temperatures below the critical temperature, $T_{\rm cr}$ [166]. This “freezing” or “condensation” phenomenon is a phase transition, since the average free energy $\langle F(T,\{E_{i}\})\rangle_{E}$ is nonanalytic at the critical temperature [168]. For Gaussian energy statistics, as in  Eq. (116), the critical temperature is $T_{\rm cr}=\sigma_{E}/\sqrt{2\ln 2}$; see  [166].
@@ 5.2.1 The 1D chains with Glauber/Metropolis dynamic ensemble
$W_{i+1,\,i}=\frac{\Gamma}{1+e^{(E_{i+1}-E_{i})/k_{B}T_{b}}}\quad\text{and}\quad W_{i-1,\,i}=\frac{\Gamma}{1+e^{(E_{i}-E_{i-1})/k_{B}T_{b}}},$ (120)
@@
$W_{i+1,\,i}=\Gamma\min\left(1,e^{-\frac{E_{i+1}-E_{i}}{k_{B}T_{b}}}\right)\quad\text{and}\quad W_{i-1,\,i}=\Gamma\min\left(1,e^{-\frac{E_{i-1}-E_{i}}{k_{B}T_{b}}}\right).$ (121)
@@
The ensemble that we have introduced can be easily studied numerically, and some limiting cases can be studied analytically. In Fig. 27, we numerically evaluate the probabilities for an even direct and inverse Mpemba index as a function of temperature, for Metropolis dynamics. Repeating the calculation for Glauber dynamics gives indistinguishable results. In both cases, the probability to have a direct Mpemba effect exceeds $0.35$ at high temperatures.
@@ 5.2.2 The random barriers ensemble
$W_{i\!j}(T_{b})=\Gamma e^{-(B_{i\!j}-E_{j})/k_{B}T_{b}},$ (122)
where $\Gamma^{-1}$ has units of time and $E_{i}$ is the energy of site $i$, and $T_{b}$ is the bath temperature. The barrier parameter $\{B_{i\!j}\}$ between sites $i$ and $j$ is defined to be symmetric, with $B_{i\!j}=B_{ji}$.
For a given system, the energies $\{E_{i}\}$ are a property of the system itself, regardless of how it is coupled to the environment or the exact nature of the environment. In contrast, the barriers, parameterized by $\{B_{i\!j}\}$, are not just internal properties of the system but rather are functions of the particularities of the environment and its precise coupling to the system. These barriers are often not known or can only be roughly estimated. In the random barrier ensemble, we treat both the energies and the barriers on an equal footing, with the barriers also randomly drawn from a distribution,
@@
where $\theta$ is the Heaviside step function. Note that to visualize the set of energies and barriers as a landscape of valleys and peaks, one would need to make sure that $B_{i\!j}\geq\{E_{i},E_{j}\}$, and while the above choice does not facilitate that, without loss of generality, one can shift all energies by a constant and ensure that, indeed, $B_{i\!j}$ exceeds $\{E_{i},E_{j}\}$.
Note that, in contrast to the ensembles discussed in Sec. (5.2.1), the system here has “mean-field coupling,” in the sense that state $i$ has non-zero transition rates to all other states. Numerical studies [96] of the probability distribution of the parity of the direct Mpemba effect show that it occurs with finite probability both above and below the freezing transition; see Fig. 28.
Figure 28: Lower bound for the direct strong Mpemba effect, showing the likelihood to have an odd number of zeros of $a_{2}$, i.e., for the probability to have a positive Mpemba index parity. The probability is evaluated as a function of the variances of the energy, $\sigma_{E}^{2}\log_{2}L$, and barrier, $\sigma_{B}^{2}\log_{2}L$. The ensemble statistics are shown for two different bath temperatures: $k_{B}T_{b}=0.1$ on panel (a), and $k_{B}T_{b}=1.0$ on panel (b), for $L=10$ energy levels. Each point was averaged over $2\times 10^{5}$ realizations. The freezing transition temperature is $T_{\rm cr}=\sigma_{E}/\sqrt{2\ln 2}\approx 0.85\,\sigma_{E}$; thus, the figures show that the probability is finite in both phases. Source: Adapted with permission from [5].
@@ 5.2.3 The isotropic ensemble
To evaluate the occurrence of the Mpemba effect across a range of systems, it is essential to determine the eigenvector corresponding to the slowest eigenvalue, the second eigenvector, for each instance in the ensemble. Solving analytically for this second eigenvector as a function of the bath temperature is challenging, particularly when dealing with large matrices. However, a method for approximating the second eigenvector was presented in [5] for a specific ensemble of large random matrices.
To understand the main idea in the isotropic ensemble, let us first consider the two ensembles discussed above in Secs. 5.2.1 and  5.2.2. In both cases, the energies are randomly drawn from a Gaussian distribution. By contrast, the barriers are either fixed by the dynamics, as in the Glauber and Metropolis cases, or chosen randomly. The eigenvectors of the relaxation dynamics are themselves random because of the randomness of the energies and, in the random-barrier case, also the barriers. In all these cases, finding the slowest eigenvector for large systems and for a given realization at arbitrary temperature can only be done numerically. In the isotropic ensemble, we avoid this problem by directly specifying the statistics of the slowest eigenvector direction rather than the barrier statistics.
To facilitate this process, it is more convenient to work with the symmetrized relaxation matrix: the relaxation rate matrix W can be transformed to a symmetric matrix $\tilde{\textbf{W}}$ via [89]
@@
where $F_{i\!j}(T_{b})=\exp(E_{i}/k_{B}T_{b})\delta_{i\!j}$. The symmetric matrix $\tilde{\textbf{W}}$ has the same eigenvalues as W and has an orthogonal set of eigenvectors $\vec{f}_{i}=F^{1/2}\vec{v}_{i}$. The eigenvectors $\{\vec{f}_{i}\}$ are orthogonal, $\vec{f}_{i}\cdot\vec{f}_{j}=||\vec{f}_{i}||^{2}\delta_{i\!j}$. Working with a symmetric matrix offers other advantages besides orthogonality; numerical diagonalization algorithms are typically more stable.
The zero-eigenvalue eigenvector of $\tilde{\textbf{W}}$ is given by $(\vec{f}_{1})_{i}\propto e^{-E_{i}/(2k_{B}T_{b})}$. It is set by choosing the energies $\{E_{i}\}$ and the bath temperature $T_{b}$. The second eigenvector of $\tilde{\textbf{W}}$ is then chosen as $\vec{g}$, a vector of i.i.d. (independent and identically distributed) random variables from a Gaussian distribution, projected onto the space orthogonal to $\vec{f}_{1}$:
@@
To estimate the occurrence of the Mpemba effect, we examine the probability associated with the parity of the Mpemba index. Below, we outline the key steps in estimating the occurrence of the direct Mpemba effect. Similarly, one can calculate the probability to have an inverse Mpemba effect.
From Eq. (42), the parity of the direct Mpemba index is
@@
where $(\vec{u}^{\,\rm dir})_{i}=\exp[E_{i}/(2k_{B}T_{b})]$ and $w_{i}=\exp[-E_{i}/(2k_{B}T_{b})]\left(\langle E\rangle_{\vec{\pi}(T_{b})}-E_{i}\right)$, and $\langle E\rangle_{\vec{\pi}(T_{b})}=\sum_{i}E_{i}\,\pi_{i}(T_{b})$ is the average energy with respect to the Boltzmann distribution at $T_{b}$. It is important to note that the vectors $\vec{u}^{\,\rm dir}$ and $\vec{w}$ depend only on the energies and not on $\vec{f}_{2}$. Because of this property, the above formula for the parity has a simple geometric meaning. Decomposing $\vec{f}_{2}$ into $\vec{f}_{||}$ and $\vec{f}_{\perp}$, components that are parallel and perpendicular to $\vec{u}^{\,\rm dir}$, we find
@@
Figure 29: Plane of components of the second eigenvector, $\vec{f}_{2}$, parallel to $\vec{u}^{\,\rm dir}$, $f_{||}$, and perpendicular to $\vec{u}^{\,\rm dir}$, $f_{\perp}$. The direct Mpemba index $\mathcal{I}^{\rm dir}_{\rm M}$ is odd in the double-wedge shaded region of the $f_{||}f_{\perp}-$plane if $\vec{u}^{\,\rm dir}\cdot\vec{w}>0$, and odd in the white regions if $\vec{u}^{\,\rm dir}\cdot\vec{w}<0$. Source: Reprinted with permission from [5].
@@
$\left(\vec{u}^{\,\rm dir}_{\rm iso}\right)_{i}\equiv e^{E_{i}/(2k_{B}T_{b})}-\frac{Le^{-E_{i}/(2k_{B}T_{b})}}{Z(T_{b},\{E_{i}\})}.$ (130)
Expressing the vector $\vec{g}$ in terms of two components, one that is parallel to $\vec{u}^{\,\rm dir}_{\rm iso}$ and denoted as $g_{||}$ and another that is orthogonal to $\vec{u}^{\,\rm dir}_{\rm iso}$ and denoted as $g_{\perp}$, we again geometrically get a double wedge where $\mathcal{P}(\mathcal{I}^{\rm dir}_{\rm M})>0$. Additionally, when estimating the probability of the Mpemba index, it is essential that the entries of $\vec{g}$ be distributed isotropically, and thus also the distribution of entries in its orthogonal projection onto $\vec{f}_{1}$. Since the components $g_{i}$ are independent and identically distributed (i.i.d.) Gaussian random variables, the same applies to $g_{||}$ and $g_{\perp}$. Consequently, we conclude that the area corresponding to the double wedge where $\mathcal{P}(\mathcal{I}^{\rm dir}_{\rm M})>0$ is defined as
@@
Prob($\mathrm{Prob}\left(\mathcal{P}(\mathcal{I}^{\rm dir}_{\rm M})>0\right)\approx\frac{C(\{E_{i}\})}{k_{B}T_{b}},$ (132)
@@
The numerical results from the random barrier ensemble are compared to the analytical estimates obtained from the isotropic ensemble in Fig. 30. In panel (a), the histograms depict the components of the second eigenvector, $(\vec{f}_{2})_{i}$, generated from the random barrier ensemble in red, alongside the components of the “candidate second eigenvector,” $(\vec{f}^{\,\,\rm{iso}}_{2})_{i}$, shown in green. Both ensembles were created for a fixed energy realization comprising $L=25$ energies drawn from a Gaussian distribution with mean zero and standard deviation $1.5$. Each ensemble consisted of $20,000$ components for the second eigenvector. The barriers were selected from a half-normal distribution with standard deviation of $75.6$. The bath temperature was set to $10$. The components of the random vectors were chosen independently and identically distributed (i.i.d.) from a Gaussian distribution with mean zero and standard deviation $0.19$. As noted in the inset, the two probability distributions exhibit differences at the tails. Figure 30b shows that the expression Eq. (131) obtained for the isotropic ensemble (solid line) captures well the probability of the direct Mpemba index for a random barrier ensemble (points) with a fixed energy realization of $L=25$ energies. This is averaged over $20,000$ randomly selected barrier realizations, based on the Eq. (123) statistics. The parameters for the statistics of energies, barriers, and the random vector $\vec{g}$ were the same as those used in Fig. 30a, although the energy sample realization was different. Both panels illustrate that the two ensembles align well when the barrier distribution and the bath temperature exceed the energy spread. In general, more work is needed to establish the range over which the average of the random barrier ensemble closely aligns with the isotropic ensemble.
@@
Figure 30: (a) Comparison of the random barrier ensemble and the isotropic ensemble for a quenched energy distribution of $L=25$ energies chosen from a Gaussian distribution with zero mean and standard deviation $\sigma_{E}=1.5$. The bath temperature is $T_{b}=10$. The sizes of the isotropic ensemble are $20,000$ random entries of $\vec{f}^{\,\,\rm iso}_{2}$. The size of the random barrier ensemble is $20,000$ random entries of $\vec{f}_{2}$. The barrier distribution is a half-normal distribution with $\sigma_{B}=75.6$; see Eq. (123). For the isotropic ensemble random vector entries, $g_{i}$ components are i.i.d.s drawn from a Gaussian with zero mean and standard deviation $\sigma_{g}=0.19$. The histogram shows the random barrier ensemble probability of $\vec{f}_{2}$ components, and the green histogram shows the isotropic ensemble histogram of the components for the “candidate second eigenvector,” $\vec{f}^{\,\,\rm iso}_{2}$. The inset emphasizes the tails of the two histograms, where we see the difference between the two ensembles. (b) Probability of the direct Mpemba index parity as a function of bath temperature $T_{b}$ for a random barrier ensemble (blue dots) and analytical estimate (orange line). The means and standard deviations of the random vector and the barriers were the same as in a). The quenched sample of $L=25$ energies was different than in a) but obeyed the same statistics (Gaussian ensemble with zero mean and standard deviation 1.5). The two ensembles match well for bath temperature and barrier distributions that are wider than the energy distribution.